# Neuron Empirical Gradient: Discovering and Quantifying Neurons’ Global Linear Controllability

### v2

@@ 8 Conclusions
This is the first study to establish a global quantitative measurement between neurons and model output, laying a foundation for precise PLM output control by modifying neurons. Our study uncovers a linear relationship between neurons in FF layers and model outputs through neuron intervention experiments. We call and quantify this linearity by “neuron empirical gradients” and propose NeurGrad, an accurate yet effective NEG estimation method. Building on NeurGrad, we deepen the understanding of neuron controllability. Finally, we demonstrate NEG’s representational ability of general language skills associated with diverse prompts through skill neuron probing experiments.
As the future work, we will investigate the neuron modification methods that can main both the preciseness and strength to apply to applications like knowledge editing and bias mitigation.
@@ 9 Limitations
Our research establishes a framework for quantitatively measuring neurons’ influence on model output and demonstrates the effectiveness of empirical gradients in representing language skills, linking language skill representation to model output through neuron empirical gradients. However, the potential for achieving skill-level model output adjustment by tuning neuron values remains unexplored. Directly adjusting neuron values could offer a more efficient alternative to traditional weight-level tuning methods. Furthermore, we also plan to examine NEG’s representational ability on language generation tasks. This approach may enable dynamic behavior modification without altering the underlying parameters of LLMs, potentially reducing computational costs and enabling more flexible model adaptation.
Furthermore, our discussion on neuron linearity and empirical gradient measurements is currently confined to single-token probing with factual prompts. In the future, we plan to expand our experiments to include prompts from diverse domains and investigate neuron attribution methods for multi-token contexts, aiming to support broader applications in generative language tasks.
@@ Appendix A Generality of Neuron Linearity
where $x$ refers to either layer or prompt, $\mathrm{maxVar}(\cdot)$ denotes the max possible variance. High coverage and distribution are desirable; a perfect generality then achieves coverage of one and distribution of one.
@@ B.1 Distribution of Neuron Activations
Figure 7: Histograms of neuron activations for six models across 1,000 prompts, displayed on a logarithmic y-axis. The figure includes three BERT models and three LLaMA-2 models, with each subplot showing the distribution of activations for one model.
In this section, we analyze the distribution of neuron activations across six PLMs, illustrated in Figure 7. The models include three BERT-based PLMs and three instruction-tuned LLaMA-2 LLMs. Figure 7 reveals that most neuron activations fall within the range of $\pm$10. While there are still some neurons that have a value out of the range of $\pm$10, the number of such neurons is comparably fewer, and increasing the range linearly increases the computational cost. Considering the balance between coverage and computational cost, we finally set the intervention range as $\pm$10 as shown in § 3.
@@ B.2 Distribution of Neuron NEGs
Figure 8: Histograms of neuron NEGs for six models across 1,000 prompts, displayed on a logarithmic y-axis.
In this section, we report the distribution of neurons’ NEG (Neuron Effect on Gradient) across the six PLMs, as illustrated in Figure 8. Similar to our discussion in § 5.1, we observe that neurons capable of altering model output are not rare. For instance, in BERTbase, over 1,000 neurons exhibit NEG magnitudes larger than 0.1. Additionally, the NEG magnitudes of neurons in LLaMA-2 LLMs are significantly smaller than those in BERT PLMs, likely due to the smaller parameter size of BERT models, which grants individual neurons greater influence over model output. Notably, all models tend to exhibit zero gradients when averaging the NEGs across all neurons.
@@ B.3 How are neurons distributed across layers?
Figure 9: Means and variances of NEG magnitudes across Transformer layers on six models. The data is calculated from the average of 1000 prompts.
We examine the variation in NEG across Transformer layers to understand the distribution of neuron controllability. Figure 9 illustrates the means and variances of magnitudes of NEG across layers. The mean NEG magnitude reflects the intensity with which PLMs adjust output probabilities through neurons in a given layer, while the variance indicates how concentrated the effective neurons are within that layer. A positive Corr is observed between variances and means, [^11] The Corr between means and variances of neuron magnitudes across different layers are 0.88, 0.79, 0.87 for BERTbase/large/wwm, and 0.57, 0.51, 0.42 for Llama2-7B/13B/70B. suggesting that as PLMs increase the intensity of gradient activity in specific layers, they also focus more on a limited subset of neurons. Specifically, in BERT models, a strong Corr is evident between layer depth and neuron controllability intensity, with deeper layers exhibiting larger gradient magnitudes, whereas Llama2 displays a distinct pattern: gradient magnitudes peak in the middle layers, decrease towards the deeper layers and then increase at the final layers. This divergence underscores the differences between the BERT and Llama2 families, emphasizing the need for case-by-case analysis in LLM mechanism investigation.
@@
(c) Evaluation results on BERTlarge.
Figure 10: Knowledge attribution evaluation by comparing CG, IG, and NeurGrad on Three BERT.
@@ Appendix C Knowledge Attribution Evaluation: Supplementary Experiments
In this section, we report the knowledge evaluation experiments on other PLMs, including three BERT PLMs. We exclude LPI from the following experiments as LPI cannot be applied to BERT models. We follow a similar experiment setup to § 4.3.
@@ Appendix D Multi-neuron Intervention: Supplementary Experiments
In this section, we report the multi-neuron intervention experiments conducted with different enhancement ranges on BERTbase and Llama2-7B, following the similar experiment setup to § 5.2. Specifically, we report the correlation between output shift and the accumulated NEGs estimated by NeurGrad with the enhancement ranges of [0, 0.1], [0, 1], [0, 1.5], and [0, 2], illustrated in Figure 11). Figure 11) demonstrates that with a larger enhancement range, involving more neurons can largely reduce the Corr, suggesting the output shift is less predictable. However, we can observe that BERTbase consistently achieves strong Corr.(>0.8) for any scenario. While Llama2-7B is less stable as BERTbase, it can still maintain moderate positive Corr. (>0.5) for 4096 neurons with enhancement range of [0,2].
@@ Appendix E Model cards
LLama2-7B(-PT) 32 11,008
Llama2-13B(-PT) 40 13,824
@@
The statistics of these six PLMs, including the number of layers (#n_layers) and neurons per layer (#neurons_per_layer) are listed in Table 5.
@@ Appendix F Construction of MCEval8K
The motivation behind creating MCEval8K is to establish a comprehensive benchmark that spans diverse knowledge genres and language skills. Since our goal is to facilitate skill neuron probing experiments where a single token must represent answers, we adopt a multi-choice task format. Additionally, we aim for the benchmark to be adaptable while avoiding redundancy for effective evaluation. In summary, we adhere to several guiding principles to design MCEval8K.
@@
    Avoid including datasets that covey similar language skills.
@@
Balanced Options: Most datasets, except for Stereoset, contain more than 8000 data points. To ensure balance across all options, we perform balanced sampling so that each option has an equal number of examples. From these datasets, we split 8000 examples into training, validation, and test sets, allocating 6,000, 1000, and 1000 examples, respectively. For instance, in the case of mLAMA, where each inquiry has four options, we ensure that the correct answer is represented equally across all four positions. This results in 1,500 occurrences (6,000/4) per position in the training set and 250 occurrences per position in both the validation and test sets.
Creation of multilingual tasks: For multilingual datasets, we focus on five languages: English (en), German (de), Spanish (es), French (fr), and Chinese (zh). These languages vary significantly in linguistic distance, with English being closer to German, French closer to Spanish, and Chinese being distant from all of them. This selection allows for a deeper analysis considering linguistic distances between languages. We ensure that 5 languages have the same number of datAn examples in each dataset (1,600 per language). Furthermore, for datasets like mLAMA, XNLI, and M-AMAZON, we ensure that each piece of knowledge is expressed in all five languages. This consistency enables direct comparisons of language understanding abilities across different languages.
@@ Appendix G Details of Skill Neuron Probing
In this section, we report the details of our skill neuron probing evaluation, including the full optimal accuracies on all tasks with zero-shot prompt setting (Table 7), few-shot prompt setting (Table 8). For two major vote probers, optimal accuracies are acquired by performing a hyper-parameter (optimal neuron size) search on the validation set and evaluating the test set. We report the optimal neuron sizes for all tasks along with the accuracies in the table.
@@
Tasks Rand LM-Prob Polar-prober (#n_neurons) Magn-prober (#n_neurons)
GED .5000 .5000 .7580 (16) .8050 (1024)
POS .2500 .5050 .5190 (16) .5470 (4)
CHUNK .2500 .3510 .4660 (8) .4490 (16)
NER .2500 .3950 .4120 (32) .4490 (8)
Agnews .2500 .4950 .6410 (32) .6900 (2)
Amazon .2000 .3750 .2750 (256) .4680 (128)
IMDB .5000 .9660 .9630 (8192) .9650 (1024)
MyriadLAMA .2500 .5080 .5200 (4) .5760 (4)
FEVER .5000 .6530 .7830 (32) .7610 (32)
CSQA .2000 .5170 .3490 (1) .5380 (16)
TempLAMA .2500 .2430 .3560 (4096) .3640 (16)
PAWS .5000 .5000 .7640 (128) .7920 (128)
MNLI .$3$3560 .4980 (4) .5590 (128)
SWAG .2500 .4610 .3360 (512) .5310 (2)
HaluEval .5000 .4990 .7540 (1024) .7510 (32)
Toxic .5000 .7230 .8250 (1024) .8210 (16)
Stereoset .$3$1096 .8299 (16) .7335 (16)
M-Amazon .2000 .2990 .2350 (4096) .3740 (2)
LTI .2000 .3670 .4300 (4) .5830 (8)
mLAMA .2500 .4020 .3880 (128) .4470 (4)
XNLI .$3$3270 .3500 (256) .3620 (16)
M-POS .2500 .3890 .2610 (1024) .3930 (4)
@@
Table 7: Optimal accuracies across all MCEval8K tasks in the zero-shot prompt setting on Llama2-7B, along with the neuron sizes achieving these accuracies.
@@
Tasks Rand LM-Prob Polar-prober (#n_neurons) Magn-prober (#n_neurons)
GED .5000 .5060 .8330 (16) .8330 (64)
POS .2500 .5730 .5870 (4) .6210 (16)
CHUNK .2500 .2710 .2820 (8192) .3910 (64)
NER .2500 .3610 .4300 (4) .4970 (64)
Agnews .2500 .5880 .7060 (64) .6890 (512)
Amazon .2000 .4840 .5310 (1) .5680 (128)
IMDB .5000 .9700 .9700 (64) .9690 (64)
MyriadLAMA .2500 .7380 .7450 (256) .7530 (4096)
FEVER .5000 .6780 .8000 (1) .8030 (4)
CSQA .2000 .6100 .6180 (32) .6340 (8192)
TempLAMA .2500 .2600 .2500 (1) .4110 (4)
PAWS .5000 .5240 .8180 (16) .8210 (32)
MNLI .$3$5100 .5780 (32) .5860 (64)
SWAG .2500 .4100 .4430 (256) .4710 (64)
HaluEval .5000 .5200 .7750 (2048) .7770 (256)
Toxic .5000 .7800 .8250 (8) .8260 (4)
Stereoset .$3$1040 .7297 (128) .5180 (16)
M-Amazon .2000 .5250 .5470 (1024) .5880 (128)
LTI .2000 .3680 .5480 (64) .6950 (8)
mLAMA .2500 .6080 .6230 (8192) .6360 (512)
XNLI .$3$3970 .4860 (32) .4980 (32)
M-POS .2500 .4440 .4830 (4) .5130 (8)
@@
Table 8: Optimal accuracies across all MCEval8K tasks in the few-shot prompt setting on Llama2-7B, along with the neuron sizes achieving these accuracies. The number of demonstrations is set as the same number of options for each task.
Appendix H Properties of Skill Neurons
H.1 Representation & Acquisition Efficiency
Neuron sizes Tasks
$2^{0}\sim 2^{3}$ Toxic, LTI, M-POS, FEVER, TempLAMA
$2^{9}\sim 2^{13}$ Agnews, MyriadLAMA, CSQA, mLAMA
Table 9: Optimal number of skill neurons in Magn-prober.
Representational efficiency: By finding the optimal neuron size on the validation set, we observe that skill-neuron prober can achieve high accuracy with a few neurons. We summarize optimal neuron sizes for all tasks with Magn-prober in Table 9. Most tasks achieved optimal accuracy within 256 neurons, demonstrating the efficiency of NEG in representing language skills. Notably, factuality tasks, such as MyriadLAMA, CSQA, and mLAMA, engage a larger number of neurons, suggesting that handling facts requires more diverse neurons, reflecting the complexity of factual understanding tasks.
Figure 12: Accuracies with varying training sizes.
Acquisition efficiency: We report the accuracy of skill-neuron probers with different training examples in Figure 12. While adding training examples can consistently increase the probers’ accuracy, the earnings slow down after 128, indicating the efficiency of acquiring skill neurons with limited data.
H.2 Generality Across Diverse Contexts
We investigate how skill neurons change when we provide different contexts, including instructions, demonstrations, and options for the same task. Given context $X$, we first acquire the skill neurons $\mathcal{N}_{s}^{X}$ and the accuracy $\text{ACC}_{\mathcal{N}_{s}^{X}}^{X}$. Then, we use the classifier built with $\mathcal{N}_{s}^{X}$ to evaluate the task by context $Y$ as $\text{ACC}_{\mathcal{N}_{s}^{X}}^{Y}$. We denote the generality of $\mathcal{N}_{s}^{X}$ on context $Y$ as $\frac{\max(\text{ACC}_{\mathcal{N}_{s}^{X}}^{Y}-\alpha,0)}{\max(\text{ACC}_{\mathcal{N}_{s}^{Y}}^{Y}-\alpha,0)}$, where $\alpha$ is the accuracy by Rand.
Using PAWS as an example, we create 12 distinct contexts by varying the instructions, the selection of demonstrations, and the output token styles. By measuring the generality for different combinations, we observe that the generality for prompting settings with different instructions and demonstrations is very high (close to 1), while the generality largely decreases if target tokens are changed. The results indicate that skill neurons maintain strong generality across different inputs, including variations in instructions and demonstrations. However, this generality diminishes when the output tokens are changed. See § K for details of experimental settings and results, including 12 designed contexts and generality results.
H.3 Are Neurons Exclusive in Skill Representations?
Figure 13: Accuracies of Magn-prober probers with different neuron sets, plotting the mean accuracy within each window, along with the accuracy ranges (min to max), as the envelope. Neuron sets are selected from all neurons in Llama2-7B in groups of 64, ranked by importance to be used as skill indicators.
We investigate whether skill neurons exclusively represent specific skills or can be substituted by different neuron sets. We thus build Magn-probers using various neuron sets. Specifically, we select 64 consecutive neurons from the ranked list, ordered by their importance as skill indicators (§ 7.1). [^12] We use 64-neurons units, which maintain high accuracies across tasks (§ A). With 352,256 neurons in Llama2-7B’s FF layers, this yields 5,504 accuracy values per task.
Figure 13 depicts the accuracies across six tasks. The result suggests skill neurons are broadly distributed, with numerous neurons acting as skill indicators. Even when relying on less important neurons, the model’s representational ability only gradually declines. Moreover, using only the least important neurons (end of each line) still yields better performance than random guesses, underscoring the inclusivity of skill neurons (See § I.3 for inclusivity evaluation on all datasets).
Appendix I Additional Analysis on Probing Results
I.1 Interpreting in-context learning with empirical gradients.
To understand why simple majority-vote classifiers achieve high accuracy, we analyze the gradients associated with each answer choice. Using PAWS (binary classification) as an example, we inspect the gradient pairs for target tokens (yes/no) across all training prompts. We find that 97.21% of neurons display opposite signs for yes/no tokens. Moreover, the Corr between yes/no gradients is -0.9996. This pronounced inverse Corr suggests that empirical gradients are sharply polarized, making it easier for a majority-vote approach to distinguish between the target tokens. Furthermore, we examine how zero-shot and few-shot prompting differ from the perspective of empirical gradients. Our analysis reveals that the total gradient magnitudes in few-shot scenarios over 22 tasks are 5.36 times greater than in zero-shot. This indicates that demonstrations in context can effectively activate skill neurons, leading to better task understanding.
I.2 More Data about Efficiency

### v3

@@ 8 Conclusions
This is the first study to provide a global quantitative link between feed-forward layer neurons and model output, and discuss the potential of precise LLM control via neuron modifications. Through neuron intervention experiments, we reveal the linear relationship between neurons and model outputs and quantify such linearity by “neuron empirical gradients” (NEG) and propose NeurGrad, an efficient NEG estimation method. Finally, we verify NEG’s ability to represent various language skills associated with diverse prompts through skill neuron probing.
As future work, we plan first to explore the trade-off between precision and intensity in neuron modification based on NeurGrad, and then leverage these insights to develop a broader neuron-level modification methodology for applications such as knowledge editing and bias mitigation.
@@ 9 Limitations
Our research establishes a framework to quantitatively measure neuron influence on model output and shows that empirical gradients effectively represent language skills, linking language skill representation to outputs via neuron empirical gradients. However, tuning neuron values for skill-level output adjustment remains unexplored and could offer a more efficient alternative to weight-level tuning. We also plan to assess NEG’s role in language generation skills, aiming for dynamic behavior changes without altering LLM parameters, reducing costs and improving flexibility in model adaptation.
Currently, our neuron linearity and gradient analysis focus on single-token factual prompts. Future work will expand to diverse domains and multi-token contexts to support wider use cases with language generation.
Acknowledgments
This paper is based on results obtained from a project, JPNP22007, commissioned by the New Energy and Industrial Technology Development Organization (NEDO).
@@ Appendix A Generality of Neuron Linearity
where $x$ refers to either layer or prompt, $\mathrm{maxVar}(\cdot)$ denotes the max possible variance. High coverage and distribution are desirable; a perfect generality then achieves coverage of one and distribution of one. The statistics in Table 5 suggest that linearity is a general property of neurons, largely independent of specific prompts or Transformer layers.
@@ B.1 Distribution of Neuron Activations
Figure 10: Histograms of neuron activations for seven models across 1,000 prompts, displayed on a logarithmic y-axis. The figure includes three BERT models and three LLaMA-2 models, with each subplot showing the distribution of activations for one model.
In this section, we analyze the distribution of neuron activations across seven PLMs, illustrated in Figure 10. The models include three BERT-based PLMs and three instruction-tuned LLaMA-2 LLMs. Figure 10 reveals that most neuron activations fall within the range of $\pm$10. While there are still some neurons that have a value out of the range of $\pm$10, the number of such neurons is comparably fewer, and increasing the range linearly increases the computational cost. Considering the balance between coverage and computational cost, we finally set the intervention range as $\pm$10 as shown in Section § 2.
@@ B.2 Distribution of Neuron NEGs
Figure 11: Histograms of neuron NEGs for seven models across 1,000 prompts, displayed on a logarithmic y-axis.
In this section, we report the distribution of neurons’ NEG across the seven PLMs, as illustrated in Figure 11. Similar to our discussion in Section § 4.1, neurons capable of altering model output are not rare. For instance, in BERTbase, over 1,000 neurons exhibit NEG magnitudes larger than 0.1. Additionally, the NEG magnitudes of neurons in LLaMA-2 LLMs are significantly smaller than those in BERT PLMs, likely due to the smaller parameter size of BERT models, which grants individual neurons greater influence over model output. Notably, all models exhibit zero gradients when averaging the NEGs across all neurons.
@@ B.3 How are neurons distributed across layers?
Figure 12: Means and variances of NEG magnitudes across Transformer layers on seven models. The data is calculated from the average of 1000 prompts.
We examine the variation in the absolute value of NEGs across Transformer layers to understand the distribution of neuron controllability. The Figure 12 illustrates the means and confidence interval of magnitudes of NEG across layers. The mean NEG magnitude reflects the intensity with which PLMs adjust output probabilities through neurons in a given layer, while the variance indicates how concentrated the effective neurons are within that layer. A positive Corr is observed between variances and means, [^14] The Corr between means and variances of neuron magnitudes across different layers are 0.88, 0.79, 0.87 for BERTbase/large/wwm, and 0.57, 0.51, 0.42 for Llama2-7B/13B/70B. suggesting that as PLMs increase the intensity of gradient activity in specific layers, they also focus more on a limited subset of neurons. Specifically, in BERT models, a strong Corr is evident between layer depth and neuron controllability intensity, with deeper layers exhibiting larger gradient magnitudes, whereas Llama2 displays a distinct pattern: gradient magnitudes peak in the middle layers, decrease towards the deeper layers and then increase at the final layers. This divergence underscores the differences between the BERT and Llama2 families, emphasizing the need for case-by-case analysis in LLM mechanism investigation.
@@
Figure 13: Knowledge attribution evaluation by comparing CG, IG, and NeurGrad on two BERT.
@@ Appendix C Knowledge Attribution Evaluation: Supplementary Experiments
In this section, we report the knowledge evaluation experiments on other PLMs, including two BERT PLMs. We exclude LPI from the following experiments as LPI cannot be applied to BERT models. We follow a similar experiment setup to Section § 3.3.
@@ Appendix D Multi-neuron Intervention: Supplementary Experiments
In this section, we report the multi-neuron intervention experiments conducted with different enhancement ranges on BERTbase and Llama2-7B, following the similar experiment setup to Section § 4.2. Specifically, we report the correlation between output shift and the accumulated NEGs estimated by NeurGrad with the enhancement ranges of [0, 0.1], [0, 1], [0, 1.5], and [0, 2], illustrated in Figure 14. Figure 14 demonstrates that a larger enhancement range, involving more neurons, can vastly reduce the Corr, suggesting the output shift is less predictable. However, we can observe that BERTbase consistently achieves strong Corr. (>0.8) for any scenario. While Llama2-7B is less stable than BERTbase, it can still maintain moderate positive Corr. (>0.5) for 4096 neurons with an enhancement range of [0,2].
@@ Appendix E Model cards
Llama3.2-3B 28 11,008
Llama2-7B 32 11,008
Qwen2.5-7B 28 18,944
Llama3.1-8B 32 14,336
@@
The statistics of these seven PLMs, including the number of layers (#n_layers) and neurons per layer (#neurons_per_layer) are listed in Table 6.
@@ Appendix F Construction of MCEval8K
The motivation behind creating MCEval8K is to establish a comprehensive benchmark that spans diverse knowledge genres and language skills. Since we aim to facilitate skill neuron probing experiments where a single token must represent answers, we adopt a multi-choice task format. Additionally, we aim for the benchmark to be adaptable while avoiding redundancy for effective evaluation. In summary, we adhere to several guiding principles to design MCEval8K.
@@
    Avoid including datasets that convey similar language skills.
@@
Balanced Options: Most datasets, except for Stereoset, contain over 8000 data points. To ensure balance across all options, we perform balanced sampling so that each option has an equal number of examples. From these datasets, we split 8000 examples into training, validation, and test sets, allocating 6,000, 1000, and 1000 examples, respectively. For instance, in the case of mLAMA, where each inquiry has four options, we ensure that the correct answer is represented equally across all four positions. This results in 1,500 occurrences (6,000/4) per position in the training set and 250 occurrences per position in both the validation and test sets.
Creation of multilingual tasks: For multilingual datasets, we focus on five languages: English (en), German (de), Spanish (es), French (fr), and Chinese (zh). These languages vary significantly in linguistic distance, with English being closer to German, French closer to Spanish, and Chinese being distant from all of them. This selection allows for a deeper analysis considering linguistic distances between languages. We ensure that 5 languages have the same number of data examples in each dataset (1,600 per language). Furthermore, for datasets like mLAMA, XNLI, and M-AMAZON, we ensure that each piece of knowledge is expressed in all five languages. This consistency enables direct comparisons of language understanding abilities across different languages.
@@ G.1 Per-task Probing Result
In this section, we report the details of our skill neuron probing evaluation, including the full optimal accuracies on all tasks with zero-shot prompt setting (Table 8) and few-shot prompt setting (Table 9). For two majority vote probes, optimal accuracies are acquired by performing a hyperparameter (optimal neuron size) search on the validation set and evaluating the test set. We report the optimal neuron sizes for all tasks and the table’s accuracies. For the random-forest probing (Tree-Probe), we directly use the gradients of all neurons to train the random forest tree. As the random forest training algorithm only takes important features to construct the decision trees, we also report the number of neurons used to build the random forests. The details of the random-forest-based probe are introduced in § G.2.
@@
Tasks Rand LM-Prob Polar-Probe (#n_neurons) Magn-Probe (#n_neurons) Tree-Probe (#n_neurons)
GED .5000 .5000 .7580 (16) .8050 (1024) 1.000 (54644)
POS .2500 .5050 .5190 (16) .5470 (4) .5850 (91290)
CHUNK .2500 .3510 .4660 (8) .4490 (16) 1.000 (93282)
NER .2500 .3950 .4120 (32) .4490 (8) 1.000 (97185)
Agnews .2500 .4950 .6410 (32) .6900 (2) .8310 (49369)
Amazon .2000 .3750 .2750 (256) .4680 (128) 1.000 (85696)
IMDB .5000 .9660 .9630 (8192) .9650 (1024) .9710 (15892)
MyriadLAMA .2500 .5080 .5200 (4) .5760 (4) 1.000 (80167)
FEVER .5000 .6530 .7830 (32) .7610 (32) .7920 (45564)
CSQA .2000 .5170 .3490 (1) .5380 (16) .5730 (96696)
TempLAMA .2500 .2430 .3560 (4096) .3640 (16) 1.000 (113786)
PAWS .5000 .5000 .7640 (128) .7920 (128) 1.000 (58200)
MNLI .$3$3560 .4980 (4) .5590 (128) .6740 (79711)
SWAG .2500 .4610 .3360 (512) .5310 (2) .5160 (96955)
HaluEval .5000 .4990 .7540 (1024) .7510 (32) 1.000 (58987)
Toxic .5000 .7230 .8250 (1024) .8210 (16) .8390 (32263)
Stereoset .$3$1096 .8299 (16) .7335 (16) .8847 (29242)
M-Amazon .2000 .2990 .2350 (4096) .3740 (2) .6260 (97623)
LTI .2000 .3670 .4300 (4) .5830 (8) .9970 (12068)
mLAMA .2500 .4020 .3880 (128) .4470 (4) .4640 (79839)
XNLI .$3$3270 .3500 (256) .3620 (16) .4510 (79212)
M-POS .2500 .3890 .2610 (1024) .3930 (4) .7740 (90001)
@@
Table 8: Optimal accuracies across all MCEval8K tasks in the zero-shot setting on Llama2-7B, with the neuron sizes achieving these accuracies.
@@
Tasks Rand LM-Prob Polar-Probe (#n_neurons) Magn-Probe (#n_neurons) Tree-Probe (#n_neurons)
GED .5000 .5060 .8330 (16) .8330 (64) 1.000 (43465)
POS .2500 .5730 .5870 (4) .6210 (16) .6550 (80695)
CHUNK .2500 .2710 .2820 (8192) .3910 (64) 1.000 (101539)
NER .2500 .3610 .4300 (4) .4970 (64) 1.000 (93577)
Agnews .2500 .5880 .7060 (64) .6890 (512) .8120 (42846)
Amazon .2000 .4840 .5310 (1) .5680 (128) 1.000 (84055)
IMDB .5000 .9700 .9700 (64) .9690 (64) .9660 (13823)
MyriadLAMA .2500 .7380 .7450 (256) .7530 (4096) .7460 (70446)
FEVER .5000 .6780 .8000 (1) .8030 (4) .8210 (38943)
CSQA .2000 .6100 .6180 (32) .6340 (8192) .6180 (94246)
TempLAMA .2500 .2600 .2500 (1) .4110 (4) 1.000 (106140)
PAWS .5000 .5240 .8180 (16) .8210 (32) 1.000 (44060)
MNLI .$3$5100 .5780 (32) .5860 (64) .6830 (67771)
SWAG .2500 .4100 .4430 (256) .4710 (64) .4160 (95311)
HaluEval .5000 .5200 .7750 (2048) .7770 (256) 1.000 (51411)
Toxic .5000 .7800 .8250 (8) .8260 (4) .8430 (29766)
Stereoset .$3$1040 .7297 (128) .5180 (16) .8204 (29774)
M-Amazon .2000 .5250 .5470 (1024) .5880 (128) .6820 (87424)
LTI .2000 .3680 .5480 (64) .6950 (8) .9910 (28362)
mLAMA .2500 .6080 .6230 (8192) .6360 (512) .6450 (75439)
XNLI .$3$3970 .4860 (32) .4980 (32) .5990 (80886)
M-POS .2500 .4440 .4830 (4) .5130 (8) 1.000 (95537)
@@
Table 9: Optimal accuracies across all MCEval8K tasks in the few-shot setting on Llama2-7B, with the neuron sizes achieving these accuracies. The number of demonstrations is set to the same number of options for each task.
G.2 Random Forest-based Probe
What is the random forest algorithm? The random forest is an ensemble learning algorithm (Heath et al., 1993) that works by creating a multitude of decision trees during training. For our multi-choice classification tasks in MCEval8K, the random forest’s output is the option selected by most trees. A decision tree is a supervised learning model that makes predictions by recursively splitting data based on feature values. During training, the tree builds nodes by selecting features that best separate the data according to a chosen metric, such as Gini impurity. Splitting continues until the data in each leaf node is sufficiently pure or a maximum depth is reached. During inference, a new input is passed through the tree by following the feature-based decisions from the root to a leaf, where the final prediction is determined by the majority label or average value of samples in that leaf.
Feature design: Our study’s objective is to explore the effectiveness of using NEG as features for knowledge representation and conduct further analysis. Therefore, the inputs for training and inference in the random forest model are constructed solely based on gradients estimated by NeurGrad. Specifically, each neuron is assigned an integer value for a given prompt. In our classification tasks, a neuron’s feature is set to $i$ if the gradient associated with the $i$-th token for that neuron is the largest among the gradients computed for all other candidate tokens (options). We ignore information on the smallest gradients to reduce the size of feature spaces.
Implementation details: For the implementation, we directly use RandomForestClassifier in scikit-learn Pedregosa et al. (2011) for training and inference. We use the default parameters of RandomForestClassifier besides the number of trees (#n_trees) and layers (#n_layers) used in each tree. The number of trees refers to the number of decision trees used to ensemble the random forest. The number of layers refers to the layer depth for each tree. Noted that RandomForestClassifier constructs binary trees; thus, the number of features used in each tree is equal to or less than $12^{\#\text{n\_layers}}-$ 1.
Visualization: We present an example of a single-tree random forest model learned from the PAWS dataset in the few-shot setting, illustrated in Figure 15. For learning this decision tree, the number of trees and layers is set to 1 and 8. The PAWS dataset is a binary classification task with candidate tokens "yes" and "no." To construct features for each neuron, we compare the NEGs computed by NeurGrad for the "yes" and "no" prompt pairs. If the gradient estimated for the prompt-yes pair exceeds that of the prompt-no pair, we assign a feature value of 1; otherwise, we assign 0.
Figure 15: Visualization of a decision tree learned for the PAWS dataset with the few-shot setting on Llama2-7B model. The “samples” in each node refers to the percentage of samples reaching this node. The “value” shows the class distribution of samples in the node.
Appendix H Additional Analysis on Probing Results
H.1 More Data about Efficiency